OpenStudy (umulas):

Need help with Partial fraction decomposition x^2 + 1 ----------------- x^3 + x^2 - 5x + 3 Which I turned into x^2 + 1 --------------- (x-1)^2(x+3)

OpenStudy (umulas):

Any ideas on what do I do now since I never did it with an addition of a numerator? Thanks,

OpenStudy (123ab456c):

=x3+x2−5x+3 EASY PEASY!

OpenStudy (umulas):

Hmmmm I would need to find them in a A = , B = , C = So it may be x^2 + 1 = A B C ------ ----- + ------ + --------- denom. x+ 3 x - 1 (x-1)^2

OpenStudy (umulas):

x^2 + 1 = A(x-1)^2 + B(x+3)(x-1) + C(x1)

OpenStudy (123ab456c):

go to math papa and press algebra calculator it will giive u the answer and explain step by step how it got the answer!:)

OpenStudy (umulas):

That's not much help, can you just double check if I'm right? I made a mistake it's x^2 + 1 = A(x-1)^2 + B(x-1)(x+3) + C (x+3)

OpenStudy (umulas):

x = 1, cancel A and B 2 = C(1+3) 2 = 4c 1/2 = C

OpenStudy (umulas):

x = -3 Cancel B and C 10 = 16A 10/16 = A 5/8 = A

OpenStudy (umulas):

x = 0 1 = 5/8 -3B + 3/2 1 = 19/8 - 3B 1 - 19/8 = -3B -11/24 = B

OpenStudy (inkyvoyd):

Okay, so you're looking to turn \(\Huge \frac{x^2 + 1 }{x^3 + x^2 - 5x + 3}\) into a partial fraction decomposition. Now, you factored, to get \(\Huge \frac{x^2 + 1 }{(x-1)^2(x+3)}\) You want to get something like this \(\Huge =\frac{A }{(x-1)^2}+\frac{B }{(x-1)}+\frac{C }{(x+3)}\) Now, to solve for the coefficients, I see you are using the heaviside cover up method. That is fine for the two non repeated terms, (x-1) and (x+3), but you will have to do the linear system for the first repeated term.

OpenStudy (inkyvoyd):

Usually in this case, since the heaviside cover up method isn't completely applicable, I'd prefer to just solve this as a system of three equation in 3 variables

OpenStudy (umulas):

5 -11 1 - ---- ---------- 8 24 2 ----- + ------- + ----------------- x + 3 x - 1 (x-1)^2

OpenStudy (inkyvoyd):

Your term for B is incorrect.

OpenStudy (umulas):

Crap, any solution for B?

OpenStudy (umulas):

9 - ---??? 24

OpenStudy (inkyvoyd):

Here, let me upload a solution.

OpenStudy (umulas):

Since 1 = 5/8 -3B + 3/2 -9/8 = -3B 9/24 = B oops it's positive

OpenStudy (umulas):

3/8

OpenStudy (inkyvoyd):

Looks good now.

OpenStudy (anonymous):

HI

OpenStudy (anonymous):

hi